OpenStudy (romantic_ch3micals):

H3LP! 1LL G1V3 4 M3D4L

OpenStudy (romantic_ch3micals):

13. Which expression is equivalent to 3c(1/3d-9)-7(c+1)+d(c+4) A)2cd-34c+4d-7 B)2cd-7c-4 C)2cd+34c+4d+7 D)2cd+8c+4 3XPL41N TOO PL34S3

OpenStudy (wmj259):

Multiply the 3c into the parenthesis. and multiply the 7 into the parenthesis. Multiply the d into the parenthesis.

OpenStudy (wmj259):

Can you draw your original question out?

OpenStudy (romantic_ch3micals):

im looking for it

OpenStudy (wmj259):

\[3c \left( \frac{ 1 }{ 3 }d-9 \right)-7\left( c+1 \right)+d \left( c+4 \right)\]

OpenStudy (wmj259):

Is that correct?

OpenStudy (romantic_ch3micals):

yea

OpenStudy (wmj259):

is the first parentheses correct?

OpenStudy (wmj259):

the d might be misplaced

OpenStudy (romantic_ch3micals):

no, its right

OpenStudy (wmj259):

okay. So use the distributive property

OpenStudy (wmj259):

To eliminate the parentheses.

OpenStudy (wmj259):

Yes?

OpenStudy (wmj259):

\[\frac{ 3c }{ 3d }-27c-7c-7+cd+4d\] Simplify

OpenStudy (wmj259):

oops, I meant to say 3/3 dc

OpenStudy (romantic_ch3micals):

.....im trying to solve it out....curse this problem

OpenStudy (wmj259):

add like terms.

OpenStudy (wmj259):

Look at my picture, take a deep breath and continue.

OpenStudy (romantic_ch3micals):

*bangs head on table*

OpenStudy (wmj259):

*No need to do that*

OpenStudy (romantic_ch3micals):

Im going for B...is it right? *drumroll*

OpenStudy (wmj259):